# Beyond Appearance: Geometric Cues for Robust Video Instance Segmentation

### v1

@@ Abstract
Corresponding author.
@@ 1 Introduction
Video Instance Segmentation (VIS) constitutes a pivotal advancement in computer vision, extending the capabilities of image-level instance segmentation to dynamic video domains [42, 29]. This task requires simultaneous segmentation, tracking, and recognition of all objects throughout video sequences, producing frame-wise outputs that include spatial positions, semantic categories, and consistent instance identities. Such capabilities underpin critical applications ranging from autonomous navigation systems to intelligent video analytics platforms [28, 2].
@@ 2 Related Work
2.1 Video Instance Segmentation
The concept of video instance segmentation (VIS) is established by [42], which first formalizes the task as joint segmentation, tracking, and recognition across video sequences. Subsequent approaches are bifurcated into online and offline paradigms: Online methods [42, 20, 40, 45, 23] sequentially process frames along the temporal dimension for real-time inference, while offline approaches [1, 39, 17, 18] leverage the full-sequence temporal context to enhance precision. With the development of transformers [34] in vision field [12, 5, 53, 10, 11], mainstream VIS methods [9, 37, 20, 40, 45, 39, 17, 18, 15, 23, 50, 51, 52, 22] turn to query-based design. Recent breakthroughs focus on decoupled designs [50, 51, 52], which delegate segmentation to query-based image models [11] and manage tracking via memory-augmented transformers. They could support both online and offline modes. Specifically, a tracker enables real-time operation in online mode, while an optional offline refiner enhances results through temporal consistency. [51] utilizing progressive denoising with contrastive learning for query propagation. [52] focuses on solving the problem of newly emerging and disappearing objects in videos using a dynamic anchor query mechanism. [22] enhances instance matching through context-aware instance features. Critically, under severe occlusions, appearance-driven methods fail to distinguish instances with homogeneous textures, leading to degraded matching reliability due to insufficient geometric constraints.
2.2 Monocular Depth Estimation
Monocular depth estimation targets reliable 3D geometric reconstruction from 2D images or videos without specialized sensors. The field has advanced toward deep learning paradigms. Early CNN-based methods [13] pioneer supervised learning using scale-invariant loss, while [14] reformulates depth prediction as an ordinal regression problem. A significant research focus has recently emerged regarding zero-shot approaches to relative monocular depth estimation. Transformative progress emerged through large-scale pretraining [32, 31, 3, 38, 19, 43, 44], where [44] notably scaled training data through synthetic image augmentation. [41] establishes a novel video monocular depth estimation framework that leverages a video generation model [4].
2.3 Depth and Segmentation Integration
@@ 3.1 Preliminary
Contemporary video segmentation predominantly adopts query-based architectures, extending image-level frameworks [10, 11] through temporal association modules for cross-frame instance matching. As illustrated in Fig. 2, our implementation follows the decoupled paradigm established in DVIS [50], consisting of: (1) a segmenter built upon the image segmentation model; (2) a tracker to propagate object associations across adjacent frames; and (3) for offline processing, a refiner applied to the tracker outputs to optimize temporal consistency throughout video sequences. The instance queries can be decoded into masks and class labels, which are the video instance segmentation results.
@@ 3.2 Expanding Depth Channel
3.3 Sharing ViT
While current state-of-the-art video segmentation methods [51, 52, 22] attain optimal performance by leveraging DINOv2  [27] pretrained Vision Transformer (ViT) [12] with ViT-Adapter [7], computational constraints often necessitate maintaining the ViT backbone in a frozen state. Meanwhile, depth estimation models inherently require backbone features for RGB-to-depth mapping. Thus, we propose a method that strategically couples the monocular depth estimation model with the segmentation model. Our design enables parameter-sharing of the powerful ViT backbone between both tasks, thereby enhancing segmentation capability while preserving depth estimation accuracy.
@@ 3.3 Sharing ViT
3.4 Depth Supervision
3.5 Implementation Details
4 Experiments
4.1 Datasets and Metrics
@@ 5 Conclusion
A promising direction involves developing enhanced depth supervision mechanisms including geometry-aware contrastive learning to eliminate inference-time dependency on depth information.

### v2

@@ Abstract
Equal contribution.Corresponding author.
@@ 1 Introduction
Video Instance Segmentation (VIS) constitutes a pivotal advancement in computer vision, extending the capabilities of image-level instance segmentation to dynamic video domains [49, 36, 28]. This task requires simultaneous segmentation, tracking, and recognition of all objects throughout video sequences, producing frame-wise outputs that include spatial positions, semantic categories, and consistent instance identities. Such capabilities underpin critical applications ranging from autonomous navigation systems to intelligent video analytics platforms [34, 2].
@@ 2 Related Work
2.1 Video Segmentation
Overview. Video Segmentation extends image segmentation to video domains, aiming to partition video frames into semantically coherent segments or object groups [28]. While preserving categorization and pixel-level segmentation capabilities, video segmentation introduces unique challenges, including temporal consistency and cross-frame object association. Following image segmentation, video segmentation tasks include video sementic segmentation (VSS), video instance segmentation (VIS) and video panoptic segmentation (VPS). Unlike three tasks mentioned above, video object segmentation (VOS) aims to track a predefined object across frames [35, 13] and reference video object segmentation (RVOS) leverages language expressions to identify the target object [23]. The latter is exemplified by [12] that utilizes motion expressions as referring cues. Recognizing inherent synergies among these tasks, emerging works [58, 27, 29] pursue unified frameworks, with [29] demonstrating a versatile architecture capable of handling both image and video segmentation even under open-vocabulary settings.
Video Instance Segmentation. The concept of video instance segmentation (VIS) is established by [49], which first formalizes the task as joint segmentation, tracking, and recognition across video sequences. Subsequent approaches are bifurcated into online and offline paradigms: Online methods [49, 22, 47, 52, 26] sequentially process frames along the temporal dimension for real-time inference, while offline approaches [1, 46, 19, 20] leverage the full-sequence temporal context to enhance precision. With the development of transformers [41] in vision field [14, 5, 60, 10, 11], mainstream VIS methods [9, 44, 22, 47, 52, 46, 19, 20, 17, 26, 57, 58, 59, 25] turn to query-based design. Recent breakthroughs focus on decoupled designs [57, 58, 59], which delegate segmentation to query-based image models [11] and manage tracking via memory-augmented transformers. They could support both online and offline modes. Specifically, a tracker enables real-time operation in online mode, while an optional offline refiner enhances results through temporal consistency. [58] utilizing progressive denoising with contrastive learning for query propagation. [59] focuses on solving the problem of newly emerging and disappearing objects in videos using a dynamic anchor query mechanism. [25] enhances instance matching through context-aware instance features. Critically, under severe occlusions, appearance-driven methods fail to distinguish instances with homogeneous textures, leading to degraded matching reliability due to insufficient geometric constraints.
2.2 Monocular Depth Estimation
Monocular depth estimation (MDE) aims to reconstruct reliable 3D geometry from 2D images or videos without specialized sensors, with recent advances predominantly driven by deep learning paradigms. Early CNN-based methods pioneered supervised learning frameworks, exemplified by [15] which introduced scale-invariant loss functions to optimize depth regression accuracy, while [16] reformulated depth prediction as an ordinal regression problem. Subsequent architectural innovation [38] uses Vision Transformer (ViT) [14] to replace convolutional networks in depth prediction tasks. A significant research shift has emerged toward zero-shot relative depth estimation, where transformative progress stems from large-scale pre-training strategies [39, 3, 45, 21, 50, 51]. [51] significantly expanded training diversity through synthetic image augmentation. Most recently, [48] establishes a novel video monocular depth estimation framework that leverages a video generation model [4].
2.3 Depth and Segmentation Integration
@@ 3.1 Preliminary
Contemporary video segmentation predominantly adopts query-based architectures, extending image-level frameworks [10, 11] through temporal association modules for cross-frame instance matching [28]. As illustrated in Fig. 2, our implementation follows the decoupled paradigm established in DVIS [57], consisting of: (1) a segmenter built upon the image segmentation model; (2) a tracker to propagate object associations across adjacent frames; and (3) for offline processing, a refiner applied to the tracker outputs to optimize temporal consistency throughout video sequences. The instance queries can be decoded into masks and class labels, which are the video instance segmentation results.
@@ 3.2 Expanding Depth Channel
3.3 Sharing ViT
While current state-of-the-art video segmentation methods [58, 59, 25] attain optimal performance by leveraging DINOv2  [33] pre-trained Vision Transformer (ViT) [14] with ViT-Adapter [7], computational constraints often necessitate maintaining the ViT backbone in a frozen state. Meanwhile, depth estimation models inherently require backbone features for RGB-to-depth mapping. Thus, we propose a method that strategically couples the monocular depth estimation model with the segmentation model. Our design enables parameter-sharing of the powerful ViT backbone between both tasks, thereby enhancing segmentation capability while preserving depth estimation accuracy.
3.4 Depth Supervision
3.5 Implementation Details
4 Experiments
4.1 Datasets and Metrics
@@ 5 Conclusion
A promising direction involves developing enhanced depth supervision mechanisms including geometry-aware contrastive learning to eliminate inference-time dependency on depth information. We hope for further investigation above our researches.